# On the Optimal Rate of Convergence for Translation-Invariant 1D Quantum Walks

### v1

@@ 3.2. Estimates on The Characteristic Functions
$\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert$$\qed$
@@
$\left\lvert\widehat{F}_{Z_{n}}^{\rho}(\lambda)-\widehat{F}_{V}^{\rho}(\lambda)\right\rvert\leq\frac{\lvert\lambda\rvert}{n}\left(\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\lvert\lambda\rvert\left(\operatorname{tr}(\lvert\mathds{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right)\right)$
@@
$\left\lvert\frac{\widehat{F}_{Z_{n}}^{\rho}(\lambda)-\widehat{F}_{V}^{\rho}(\lambda)}{\lambda}\right\rvert$ $\leq n^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\lvert\lambda\rvert\left(\operatorname{tr}(\lvert\mathds{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right)\right)$
$\leq n^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert\mathds{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).$
@@
$\left\lvert\frac{\widehat{F}_{Z_{n}}^{\rho}(\lambda)-\widehat{F}_{V}^{\rho}(\lambda)}{\lambda^{2}}\right\rvert$ $\leq n^{-1}\left(\lvert\lambda\rvert^{-1}\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert\mathds{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right)$
$\leq n^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert\mathds{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).$
@@
$L(\widehat{F}_{Z_{n}}^{\rho}-\widehat{F}_{V}^{\rho})$ $\leq\epsilon+\epsilon^{-2}Cn^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert\mathds{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).$
@@
$L(\widehat{F}_{Z_{n}}^{\rho}-\widehat{F}_{V}^{\rho})$ $\leq n^{-\frac{1}{3}}\left(1+\sum_{k\in{\mathcal{I}}}\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert\mathds{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).\qed$
@@ 4. Berry-Esseen Bounds for Coin-Step Walks
In this Section, we prove Theorem˜2.2. We thus now restrict our attention to coin-step walks with two degrees of freedom as introduced in Section˜2.3 with the general unitary coin matrix $C$ given by
@@ 4.1. The Asymptotic Distribution
$\sup_{\lvert x\rvert\leq\lvert a\rvert-r}\left\lvert F_{{\mathds{X}}_{n}}^{\phi}(x)-F_{V}^{\phi}(x)\right\rvert\leq C_{r}n^{-\frac{1}{3}}.$
@@
For the wavefront region, our results heavily rely on the explicit expression for the density of $F_{{\mathds{X}}_{n}}^{\phi}$, calculated by Sunada and Tate in [ST12]. In the following, we recall and extend some of their results. We then apply them to prove lower and upper bounds on $F_{{\mathds{X}}_{n}}^{\phi}$ in close proximity to $\pm|a|$. We remark that, further outside the propagation region, exponential decay of occupation probability has been proven [ST12, CJWW25].
@@
(4.1) $p_{n}(\phi;k)\coloneqq\operatorname{tr}\big(\rho\mathbf{1}_{\{k\}}W^{-n}{\mathds{X}}W^{n}\big),\qquad F_{n}(\phi;x)\coloneqq\sum_{k\leq x}p_{n}(\phi;k),$
@@
$F_{V}^{\phi}(-|a|+n^{-\frac{2}{3}-\varepsilon})+n^{-\frac{2}{3}-\varepsilon}\lesssim n^{-\frac{1}{3}-\varepsilon/2}+n^{-\frac{2}{3}-\varepsilon}\lesssim n^{-\frac{1}{3}}\lesssim F_{{\mathds{X}}_{n}}^{\phi}(-|a|),$
which proves $n^{-2/3-\varepsilon}\lesssim L(F_{V}^{\phi},F_{{\mathds{X}}_{n}}^{\phi})$. ∎
@@
For the purpose of this proof, we consider $F_{n}(\phi;\cdot)$ on the three regions $(-\infty,-Cn^{\frac{2}{3}}],[-Cn^{\frac{2}{3}},-n^{\frac{1}{3}}],[-n^{\frac{1}{3}},n^{\frac{1}{3}}]$, where we pick $C>0$ such that $L(F_{{\mathds{X}}_{n}}^{\phi},F_{V}^{\phi})<Cn^{-1/3}$ for all $n\in{\mathbb{N}}$. This is possible by Theorem˜2.1.
For the first region, we directly apply the bound on the Lévy metric Theorem˜2.1 and the fact that $F_{V}^{\phi}(k)=0$ for $k\leq-|a|$, cf. Proposition˜4.1. With the choice $L_{n}\coloneqq\frac{1}{2}L(F_{{\mathds{X}}_{n}}^{\phi},F_{V}^{\phi})+\frac{1}{2}Cn^{-1/3}=O(n^{-1/3})$, this yields
$F_{n}(\phi;-n\lvert a\rvert-Cn^{\frac{2}{3}})=F_{{\mathds{X}}_{n}}^{\phi}(-|a|-Cn^{-\frac{1}{3}})\leq F_{V}^{\phi}(-\lvert a\rvert-Cn^{-\frac{1}{3}}+L_{n})+L_{n}\leq 0+Cn^{-\frac{1}{3}}.$
@@
$\textup{OSC}^{\pm}_{n}\left(\tfrac{k}{n}\right)\coloneqq p_{n}\left(\phi,\pm\lfloor n\lvert a\rvert\rfloor\mp k\right)-\frac{1}{n}\sigma_{C,\phi}\left(\tfrac{k}{n}\right)$
@@
$\qquad\leq\left\lvert F_{{\mathds{X}}_{n}}^{\phi}(\lvert a\rvert+r^{\prime})-F_{V}^{\phi}(-\lvert a\rvert+r^{\prime})\right\rvert+O(n^{-\frac{1}{3}})\xrightarrow{n\to\infty}0,$

### v2

@@ 3.2. Estimates on The Characteristic Functions
$\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert$$\qed$
@@
$\left\lvert\widehat{F}_{Z_{n}}^{\rho}(\lambda)-\widehat{F}_{V}^{\rho}(\lambda)\right\rvert\leq\frac{\lvert\lambda\rvert}{n}\left(\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\lvert\lambda\rvert\left(\operatorname{tr}(\lvert{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right)\right)$
@@
$\left\lvert\frac{\widehat{F}_{Z_{n}}^{\rho}(\lambda)-\widehat{F}_{V}^{\rho}(\lambda)}{\lambda}\right\rvert$ $\leq n^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\lvert\lambda\rvert\left(\operatorname{tr}(\lvert{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right)\right)$
$\leq n^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).$
@@
$\left\lvert\frac{\widehat{F}_{Z_{n}}^{\rho}(\lambda)-\widehat{F}_{V}^{\rho}(\lambda)}{\lambda^{2}}\right\rvert$ $\leq n^{-1}\left(\lvert\lambda\rvert^{-1}\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right)$
$\leq n^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).$
@@
$L(\widehat{F}_{Z_{n}}^{\rho}-\widehat{F}_{V}^{\rho})$ $\leq\epsilon+\epsilon^{-2}Cn^{-1}\left(\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).$
@@
$L(\widehat{F}_{Z_{n}}^{\rho}-\widehat{F}_{V}^{\rho})$ $\leq n^{-\frac{1}{3}}C\left(C^{-1}+\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\operatorname{tr}(\lvert{X}\rvert\rho)+\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{\mathcal{B}(\mathcal{H})}\right).\qed$
@@ 4. Berry-Esseen Bounds for Coin-Step Walks
In this Section, we prove Theorem˜2.2. We thus now restrict our attention to shift-coin walks with two degrees of freedom as introduced in Section˜2.3 with the general unitary coin matrix $C$ given by
@@ 4.1. The Asymptotic Distribution
$\sup_{\lvert x\rvert\leq\lvert a\rvert-r}\left\lvert F_{X_{n}}^{\phi}(x)-F_{V}^{\phi}(x)\right\rvert\leq C_{r}n^{-\frac{1}{3}}.$
@@
For the wavefront region, our results heavily rely on the explicit expression for the density of $F_{X_{n}}^{\phi}$, calculated by Sunada and Tate in [ST12]. In the following, we recall and extend some of their results. We then apply them to prove lower and upper bounds on $F_{X_{n}}^{\phi}$ in close proximity to $\pm|a|$. We remark that, further outside the propagation region, exponential decay of occupation probability has been proven [ST12, CJWW25].
@@
(4.1) $p_{n}(\phi;k)\coloneqq\operatorname{tr}\big(\rho\mathbf{1}_{\{k\}}W^{-n}XW^{n}\big),\qquad F_{n}(\phi;x)\coloneqq\sum_{k\leq x}p_{n}(\phi;k),$
@@
$F_{V}^{\phi}(-|a|+n^{-\frac{2}{3}-\varepsilon})+n^{-\frac{2}{3}-\varepsilon}\lesssim n^{-\frac{1}{3}-\varepsilon/2}+n^{-\frac{2}{3}-\varepsilon}\lesssim n^{-\frac{1}{3}}\lesssim F_{X_{n}}^{\phi}(-|a|),$
which proves $n^{-2/3-\varepsilon}\lesssim L(F_{V}^{\phi},F_{X_{n}}^{\phi})$. ∎
@@
For the purpose of this proof, we consider $F_{n}(\phi;\cdot)$ on the three regions $(-\infty,-Cn^{\frac{2}{3}}],[-Cn^{\frac{2}{3}},-n^{\frac{1}{3}}],[-n^{\frac{1}{3}},n^{\frac{1}{3}}]$, where we pick $C>0$ such that $L(F_{X_{n}}^{\phi},F_{V}^{\phi})<Cn^{-1/3}$ for all $n\in{\mathbb{N}}$. This is possible by Theorem˜2.1.
For the first region, we directly apply the bound on the Lévy metric Theorem˜2.1 and the fact that $F_{V}^{\phi}(k)=0$ for $k\leq-|a|$, cf. Proposition˜4.1. With the choice $L_{n}\coloneqq\frac{1}{2}L(F_{X_{n}}^{\phi},F_{V}^{\phi})+\frac{1}{2}Cn^{-1/3}=O(n^{-1/3})$, this yields
$F_{n}(\phi;-n\lvert a\rvert-Cn^{\frac{2}{3}})=F_{X_{n}}^{\phi}(-|a|-Cn^{-\frac{1}{3}})\leq F_{V}^{\phi}(-\lvert a\rvert-Cn^{-\frac{1}{3}}+L_{n})+L_{n}\leq 0+Cn^{-\frac{1}{3}}.$
@@
$\textup{OSC}^{\pm}_{n}\left(\tfrac{k}{n}\right)\coloneqq p_{n}\left(\phi,\pm\lfloor n\lvert a\rvert\rfloor\mp k\right)-\frac{1}{n}\sigma_{C,\phi}\left({\pm\lvert a\rvert\mp}\tfrac{k}{n}\right)$
@@
$\qquad\leq\left\lvert F_{X_{n}}^{\phi}(\lvert a\rvert+r^{\prime})-F_{V}^{\phi}(-\lvert a\rvert+r^{\prime})\right\rvert+O(n^{-\frac{1}{3}})\xrightarrow{n\to\infty}0,$